OpenStudy (anonymous):

3. Find the fifth term in this arithmetic sequence: 3, 7,

OpenStudy (anonymous):

9,11,13.

OpenStudy (lgbasallote):

you have A1 = 3 you're looking for An which is A5 in this problem.. n = 5 because a_n = a_5 d = a_2 - a_1 = 7 - 3 = 4 An = A1 + (n-1)d A5 = 3 + (5-1)4